Sol Summary – November 29th

Sol 2

Title: An Intriguing Find…

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, I set off to explore the area to the South of Kissing Camel Range and West of Cow Dung Rd. First I investigated some of the washes crossing the road to see what I could find. In the first wash that I stopped in, I found pieces of petrified wood amongst other very interesting specimens. I’ll definitely have to explore this drainage basin later.

From there, I proceeded further south, until I reached a large wash near White Rock Canyon. I only walked about 2 minutes up, when I found a very large specimen that looked like mineralized bone marrow. Compared to the basalt ejecta littering the wash, this specimen looked completely different. The holes were all aligned in the same direction, and were about the same size. The basalt pieces had gas bubbles, but were of random size, with no obvious pattern. The size of the marrow was very large, with 3-4 mm spacing. If this was indeed a piece of a dinosaur bone, then that meant there is a very large dinosaur fossil somewhere in that drainage basin.

I left the specimen where I found it, planning to come back and visit later. After spending over 2 hours in this area, I didn’t find much more, though there is still quite a bit of ground to cover. I returned to the hab by around 2:00 PM, and informed Shannon on what I had found. On our way to town, we investigated the specimen that I had found in the wash. Shannon wasn’t sure what it was, but we confirmed that there weren’t any other rocks in that area that looked like it.  She figured that if it was indeed a remnant of a dinosaur bone, that it would have to come from a very large dinosaur, much larger than a Utahraptor or velociraptor. We plan on surveying this area further to see what we can find.

Soon after, we ran into town, found a socket wrench kit that would work with the bolts on the new rovers, and organized the next supply of food for Crew 184. We returned to hab after finishing our business in town.

Look Ahead Plan:

Tomorrow and the day after, I will be extensively searching the drainage basin where I found that specimen for large fossil exposures. Since Shannon told me that she hasn’t extensively explored this area, what I may find there could be very interesting.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy


Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – November 29th

Crew 183 + Operations Report 29NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 5:30 pm

Solar— SOC  91% (Before generator is run at night)

Diesel –  50 %

Propane –  100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  no


Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:


Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:  Propane was delivered today.  A leak was
found in the line of the small tank at the director’s house, and it
was fixed.  We began work on the second rover today but then decided
to go to storage and see it we could find better tools.  We found a
socket set that was the right size!  We will get back on it tomorrow.
We also brought a 55 gallon drum from town.  Tomorrow I will pound
holes in it and set it up as a burn barrel.  I will also need to clean
up the old burn barrel area as non-burnable garbage was burned and it
will need to be picked up and taken for disposal.  Lucky me!

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Three new smoke detectors arrived today.

Summary of GreenHab operations: Door was opened this afternoon.  Two
gallons of water was used in the new planters.  I did not water the
plants today.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Thanks for all of
the help with the rover installation question last night.  No
questions tonight.

Copyright © The Mars Society. All rights reserved. | Main Site